### v1

# Solving Sequential Greedy Problems Distributedly with Sub-Logarithmic Energy Cost

Abstract
We study the awake complexity of graph problems that belong to the class O-LOCAL, which includes a large subset of problems solvable by sequential greedy algorithms, such as $(\Delta+1)$-coloring, maximal independent set, maximal matching, etc. It is known from previous work that, in $n$-node graphs of maximum degree $\Delta$, any problem in the class O-LOCAL can be solved by a deterministic distributed algorithm with awake complexity $O(\log\Delta+\log^{\star}n)$.
In this paper, we show that any problem belonging to the class O-LOCAL can be solved by a deterministic distributed algorithm with awake complexity $O(\sqrt{\log n}\cdot\log^{\star}n)$. This leads to a polynomial improvement over the state of the art when $\Delta\gg 2^{\sqrt{\log n}}$, e.g., $\Delta=n^{\epsilon}$ for some arbitrarily small $\epsilon>0$. The key ingredient for achieving our results is the computation of a network decomposition that uses a small-enough number of colors in sub-logarithmic time in the Sleeping model, which can be of independent interest.
Keywords: distributed graph algorithms, energy-efficient algorithms, Sleeping model.
1 Introduction
In the last four decades, the study of the distributed complexity of graph problems has mainly been focused on understanding the worst-case complexity. Recently, however, different notions of complexities, that better capture the energy usage of a distributed system, have attracted lots of attention. One successful such notion of complexity is captured by the Sleeping variant of the LOCAL model.
The Sleeping model has been introduced in [CGP20]. It assumes a set of $n$ fault-free processes connected by an $n$-node graph $G=(V,E)$, in which computation proceeds in lockstep, as a sequence of synchronous rounds. At each round, each process, i.e., each node of $G$, is either awake or asleep. A node $v\in V$ that is awake at a round $r$ can send a message (of arbitrary size) to each of its neighbors, receive the messages sent by its (awake) neighbors, and perform some individual computation (the local computation power of each node is not bounded). Every node $v$ has also the ability to become asleep for a prescribed number of rounds, whose value is chosen by $v$ as a function of its internal state. If node $v$ remains awake at the end of round $r\geq 1$, then it directly proceeds to round $r+1$. Otherwise, it becomes asleep for, say $t\geq 1$ rounds (where $t$ is chosen by $v$). A sleeping node cannot send messages, and all messages sent to a sleeping node are lost. After $t$ rounds, node $v$ wakes up, at round $r+t+1$. Initially, i.e., at round 1, all nodes are awake. The main motivation for the study of the Sleeping model is to determine the power and limitation of distributed systems in which processing nodes have the ability to save energy by turning themselves off for a prescribed amount of time.
In addition to the classical round complexity measuring the number of rounds required to solve a given problem, the main measure of complexity in the Sleeping model is the awake complexity, which is the maximum, taken over all the nodes $v\in V$, of the number of rounds during which $v$ is awake.
This paper focuses on the awake complexity of solving a large class of natural problems, namely those in the class O-LOCAL, introduced in [BM21]. Roughly, this class includes all problems which can be solved by a restricted form of sequential greedy algorithms. More specifically, the order in which the nodes are sequentially picked is governed by an arbitrary given acyclic orientation of the edges of the graph $G$, that is, a node $v$ can be picked only if solutions have been computed for all its descendants according to the given orientation. More importantly, the solution at $v$ can be computed based only on the solutions previously computed for these descendants. Even if O-LOCAL does not contain all sequentially greedily solvable graph problems (e.g., distance-2 coloring is not in O-LOCAL), it includes many among the most important problems in distributed network computing, such as, $(\Delta+1)$-vertex coloring, maximal independent set (MIS), maximal matching, $(2\Delta-1)$-edge coloring, etc, where $\Delta$ denotes the maximum degree in the graph.
[…]
1.1 Our Results
We exhibit a quadratic improvement for the awake complexity of O-LOCAL problems, when expressed as a function of $n$. Specifically, we show the following result.
Theorem 1.
[…]
In particular, for $\Delta\gg 2^{\sqrt{\log n}\cdot\log^{\star}n}$, e.g., $\Delta=n^{\epsilon}$ for some arbitrarily small $\epsilon>0$, this improves the algorithm from [BM21]. The round complexity of our algorithm is $O(n^{5}\sqrt{\log n})$. However, if the IDs assigned to the $n$ nodes are taken from $\{1,\ldots,n\}$, then this round complexity improves to $O(n^{2}\sqrt{\log n})$. More generally, if the IDs are taken from $\{1,\ldots,n^{c}\}$ with $c\geq 1$, then the round complexity of our algorithm becomes $O(n^{\min\{1+c,5\}}\sqrt{\log n})$.
Our result is obtained thanks to an interplay between two similar forms of network decomposition. Both are partitioning the nodes of the input graph $G=(V,E)$ into clusters (i.e., connected induced subgraphs of $G$). Every cluster $C$ has a root $r$, and a label $x$. It is encoded distributedly by having each node $v\in C$ storing a pair $(\ell(v),\delta(v))$ where $\ell(v)=x$, and $\delta(v)$ is the distance from $v$ to $r$ in $C$. We thus refer to our network decomposition as BFS-clustering. Our two forms of clustering differ according to the labeling of the clusters. The uniquely-labeled BFS-clustering requires that no two clusters have the same label. Instead, the colored BFS-clustering allows different clusters to have the same label, as long as they are not adjacent. That is, clusters $C$ and $C^{\prime}$ of $G=(V,E)$ may have the same label, but then it must be the case that, for every $v\in C$ and every $v^{\prime}\in C^{\prime}$, $\{v,v^{\prime}\}\notin E$.
The former type of clustering enables induction. In fact, in the virtual graph resulting from merging each cluster into a single vertex, it is possible to run algorithms for the LOCAL model with only a constant overhead on the awake complexity. The latter type of clustering is weaker, yet it is sufficient to apply previous results of the literature on that form of clustering, such as the one in [BM21], which assumes a given proper $k$-coloring.
[…]
1.2 Additional Related Work
The Sleeping model was introduced in [CGP20], which opened a large avenue of research dedicated to measuring the potential benefit of providing the nodes of a distributed system with the ability to turn on and off at will. In particular, MIS has attracted a lot of attention in this context (see [CGP20, GP22, GP23, HPR22]), and it was eventually shown [DJP23] that MIS can be solved (by a randomized algorithm) with awake complexity $O(\log\log n)$. Regarding $(\Delta+1)$-coloring, the aforementioned paper [BM21] on solving O-LOCAL problems established a (deterministic) awake complexity $O(\log\Delta+\log^{\star}n)$. The randomized algorithm in [DFRZ24] has awake complexity $O(\log\log n)$. This presents an improvement over the best known randomized algorithms for $(\Delta+1)$-coloring in the LOCAL model [CLP18, HKNT22], which have round complexity $\tilde{O}(\log^{2}\log n)$, though to the expense of a significant increase of the round complexity. Several other problems have been studied in the Sleeping model, including matching [GP22], vertex cover [GP22], spanning tree [BM21], and minimum-weight spanning tree [AJP24]. Note that problems such as maximal matching and $(2\Delta-1)$-edge-coloring are in O-LOCAL, and can thus be solved with awake complexity $O(\log\Delta+\log^{\star}n)$ [BM21]. A direct consequence of the result in this paper is that MIS, $(\Delta+1)$-coloring, maximal matching, and $(2\Delta-1)$-edge-coloring have awake complexity $O(\sqrt{\log n}\cdot\log^{\star}n)$.
We note that notions that are closely related to the one of awake complexity have been investigated in the context of radio networks (see e.g., [JKZ02b, CKP+19, NO00, KKP13, JKZ02a, CDHP20, BKPY18]), and also in the context of the LOCAL model, such as the line of work that studies the average complexity of graph problems in the LOCAL model (see e.g., [BT19, Feu20, BBK+24, BBK+23, BGKO22]).
2 Model and Definitions
2.1 The Sleeping Model
We consider the Sleeping variant of the standard LOCAL model [Pel00]. Recall that the LOCAL model assumes $n$ nodes connected as an $n$-node network, modeled as a simple connected graph $G=(V,E)$. Every node has an identifier, which is unique in the network. The identifiers are supposed to be taken from a polynomial range of identifiers, and thus every identifier can be stored on $O(\log n)$ bits in $n$-node graphs. Computation proceeds synchronously as a sequence of rounds. All nodes starts at round 1. At each round, every node can send a message to each of its neighbors, receive the messages from its neighbors, and perform some internal computation. There are no limits on the size of the messages, nor on the amount of internal computation a node may perform at each round.
At each round of a distributed algorithm for the LOCAL model, the Sleeping model provides each node $v$ with the additional capacity to become “asleep” (i.e., to turn off) for a prescribed number of rounds, which is chosen at will by $v$. A node executing an algorithm in the Sleeping model thus alternates between sequences of rounds during which it is “awake” (i.e., is on), and sequences of rounds during which it is asleep. If a node is awake at a given round, it acts like in the LOCAL model. However, a node $v$ that is asleep at a round is totally inactive at this round. In particular, all messages that may have been sent to $v$ by its neighbors at this round are lost, and they cannot be recovered (unless sent again by the neighbors at some later round during which $v$ is awake). Initially, all nodes are awake. We assume that every processing node $v$ initially knows the order $n$ of the graph $G=(V,E)$ it belongs to.
The round complexity of an algorithm $A$ running under the Sleeping model is the maximum, taken over all graphs $G$ of at most $n$ nodes, and over all nodes $v$ of $G$, of the number of rounds required for $v$ to terminate. The awake complexity of $A$, is the maximum, taken over all graphs $G$ of at most $n$ nodes, and over all nodes $v$ of $G$, of the number of rounds during which $v$ is awake.
2.2 The O-LOCAL Class of Graph Problems
O-LOCAL, which stands for Oriented-LOCAL [BM21], is a class of graph problems whose every solution consists of an appropriate labeling of the nodes of the input graph (e.g., $(\Delta+1)$-coloring, maximal independent set (MIS), etc.) that can be computed by a sequential greedy algorithm performing as follows. Let $G=(V,E)$ be a (non-directed) graph, let $\mu$ be an arbitrary acyclic orientation of the edges of $G$, and let $G_{\mu}$ be the resulting directed graph. For every $v\in V$, let $G_{\mu}(v)$ be the subgraph of $G_{\mu}$ induced by $v$ and all the nodes that can be reached from $v$ by following outgoing edges only. That is, $G_{\mu}(v)$ contains all nodes $w$ such that there is a path from $v$ to $w$ in the directed graph $G_{\mu}$. The greedy algorithm picks nodes in arbitrary order, but respecting the orientation $\mu$. That is, a node $v$ can be picked only if the output of all the nodes in $G_{\mu}(v)\smallsetminus\{v\}$ has been previously computed by the algorithm. Let $v$ be such a node. The algorithm must be able to compute a correct output for $v$ based only on the outputs computed previously by all nodes in $G_{\mu}(v)\setminus\{v\}$. A graph problem $\Pi$ is in O-LOCAL if there is a greedy algorithm that succeeds to compute a solution for $\Pi$ for every input graph $G$, and for every acyclic orientation $\mu$ of the edges of $G$.
The class O-LOCAL is, by definition, a subset of all graph problems that can be solved by a greedy algorithm fixing the outputs of the nodes by treating them sequentially, in arbitrary order. In particular, O-LOCAL contains problems at the core of distributed computing in networks, including (proper) $(\Delta+1)$-coloring of graphs with maximum degree $\Delta$, and maximal independent set (MIS). Distance-2 $(\Delta^{2}+1)$-coloring is however not in O-LOCAL, as witnessed by the $n$-node path $P$, with $n\geq 6$. Indeed, let $\mu$ be the edge orientation of $P$ such that every two incident edges are directed in opposite direction. For this orientation, the color of every node $v$ with out-degree 0 in $P_{\mu}$ must be fixed without any other information than the ID of $v$. For every $f:\{1,\dots,n\}\to\{1,\dots,5\}$, there is an assignment of IDs to the nodes of $P$ such that the color assignment $f$ collides for two vertices at distance 2 in $P$.
2.3 Definitions
We define a collection of concepts that will be at the core of the technical parts of the paper. Specifically, we define two forms of network decomposition, a.k.a. clustering, and we associate a virtual graph to each of these two decompositions, referred to as uniquely-labeled BFS-clustering and colored BFS-clustering. BFS refers to the fact that, in each cluster, there is a special node, called the root, and every node in the cluster knows its distance to the root. A uniquely-labeled BFS-clustering assigns a label to each cluster, and this label is unique, i.e., no two clusters have the same label. This enables to recurse in the virtual graph resulting from collapsing every cluster into a single vertex, and assigning the label of the cluster as ID of the vertex. If one enforces conditions on the number of labels, computing a uniquely-labeled BFS-clustering is not easy. In such a case, colored BFS-clusterings are easier to compute, but two distinct cluster may be given the same label (now refer to as color), as long as there are no edges between them in the graph. The issue with colored BFS-clusterings is that one may not be able to recurse in the virtual graphs induced by these BFS-clustering, whenever considering algorithms in the LOCAL model, which assume distinct IDs, yet they are sufficient for applying known results (e.g., algorithms in [BM21], which assume a given $k$-coloring). We thus play with the two types of BFS-clustering for benefiting of the best of each of them.
Definition 2.
A uniquely-labeled BFS-clustering of a graph $G=(V,E)$ is a pair of functions $(\ell,\delta)$ assigning a pair $(\ell(v),\delta(v))\in\mathbb{N}\times\mathbb{N}$ to each node $v\in V$ such that, for every integer $i>0$, if the subgraph $G_{i}$ of $G$ induced by the nodes in $\{v\in V\mid\ell(v)=i\}$ is non-empty, then $G_{i}$ is connected, there is a unique node $u$ of $G_{i}$ with $\delta(u)=0$, and, for every node $v$ of $G_{i}$, $\delta(v)$ is the distance from $u$ to $v$ in $G_{i}$.
Given a uniquely-labeled BFS-clustering $(\ell,\delta)$ of $G=(V,E)$, and $i\in\mathbb{N}$, the set of nodes $\{v\in V\mid\ell(v)=i\}$ is called a cluster, $\ell(v)$ is called the cluster-ID of $v\in V$, and we somewhat abuse terminology by saying that node $v$ with $\ell(v)=i$ belongs to the cluster $i$, or the $i$-th cluster.
Definition 3.
The virtual graph induced by a uniquely-labeled BFS-clustering $(\ell,\delta)$ of a graph $G=(V,E)$ is the graph $H$ with vertex-set equal to the set $\{\ell(v)\mid v\in V\}$ of cluster-IDs, and two vertices $i$ and $j$ of $H$ are neighbors in $H$ if there exists an edge $\{u,v\}\in E$ such that $\ell(u)=i$ and $\ell(v)=j$.
That is, $H$ can be viewed as the graph obtained from $G$ by merging each cluster of the uniquely-labeled BFS-clustering into one vertex, and replacing every set of parallel edges resulting from this merge by a single edge.
Definition 4.
A colored BFS-clustering of a graph $G=(V,E)$ is a pair of functions $(\gamma,\delta)$ assigning a pair $(\gamma(v),\delta(v))\in\mathbb{N}\times\mathbb{N}$ to every node $v\in V$ such that, for every $i\in\mathbb{N}$, the subgraph $G_{i}$ of $G$ induced by the nodes $\{v\in V\mid\gamma(v)=i\}$ satisfies that, for every connected component $C$ of $G_{i}$, there is a unique node $u$ of $C$ with $\delta(u)=0$, and, for every $v\in C$, $\delta(v)$ is the distance between $u$ and $v$ in $C$.
Given a colored BFS-clustering $(\gamma,\delta)$ of a graph $G=(V,E)$, for every $i\in\mathbb{N}$, and every connected component of the subgraph $G_{i}$ of $G$ induced by the nodes $\{v\in V\mid\gamma(v)=i\}$, we say that $C$ is a cluster, and that $i$ is the color of that cluster. Note that, in a uniquely-labeled BFS-clustering of a graph $G$, each set of nodes with identical label is connected, and the ID of the corresponding cluster is unique (no other clusters have the same ID). In contrast, in a colored BFS-clustering, non-adjacent clusters are allowed to have the same ID, which motivates the terminology “color”.
Definition 5.
The virtual graph induced by a colored BFS-clustering of a graph $G=(V,E)$ is the graph $H$ with vertex-set equal to the set of clusters resulting from the clustering, and two vertices $C$ and $C^{\prime}$ of $H$ are neighbors in $H$ if there exist $\{u,v\}\in E$ such that $u\in C$ and $v\in C^{\prime}$.
As for uniquely-labeled BFS-clustering, the virtual graph $H$ induced by a colored BFS-clustering can be viewed as the graph obtained from $G$ by merging each cluster $C$ of the clustering into one vertex, and replacing every set of parallel edges resulting from this merge by a single edge.
3 Prelimilaries
In this section, we prove some useful statements about the Sleeping model. We start by stating the following lemma, which is implied by Lemma 2.1 in [BM21], but for which we provide a sketch of proof that can serve as a warm up for the reader unfamiliar with the Sleeping model.
Lemma 6 (Barenboim and Maimon [BM21]).
Let $G=(V,E)$ be a graph, and let $T$ be a spanning tree of $G$ rooted at $r\in V$. For every $v\in V\smallsetminus\{r\}$, let $p(v)$ denote the parent of $v$ in $T$. Let $L:V\to\{1,\dots,N\}$ be a labeling such that $L(v)>L(p(v))$ for every $v\in V\smallsetminus\{r\}$. Let us assume that every node $v$ is given its parent $p(v)$ (if $v\neq r$), its label $L(v)$, and $N$. Each of the two tasks broadcast and convergecast can be performed in $G$ under the Sleeping model, with awake complexity 3, and round complexity $O(N)$.
Sketch of proof.
Let $M$ be the message to be broadcasted by $r$ to all the nodes. At the first round, every node $v$ communicates with its neighbors for learning the label $L(p(v))$ of its parent, and becomes idle. Every node then wakes up only twice. Specifically, every node $v$ wakes up at round $2+L(p(v))$ during which it communicates with its neighbors for learning $M$ (that it receives from its parent). Node $v$ then becomes idle, and wakes up again at round $2+L(v)$ during which it sends $M$ to all its neighbors, and thus in particular to its children in $T$.
Convergecast is performed in a way similar to broadcast. For every node $v\in V$, let $M_{v}$ be the message of $v$. First, every node $v$ assigns itself a new label $L^{\prime}(v)=N-L(v)$. Note that this new labeling satisfies $L^{\prime}(v)<L^{\prime}(p(v))$ for every $v\neq r$. At the first round, every node $v$ communicate with its neighbors for learning the label $L^{\prime}(p(v))$ of its parent, and becomes idle. As for broadcast, every node then wakes up only twice. Specifically, node $v$ wakes up at round $2+L^{\prime}(v)$ during which it receives from the children all convergecast messages from nodes in the subtrees of $T$ rooted at these children. Node $v$ then becomes idle, and wakes up again at round $2+L^{\prime}(p(v))$ during which $v$ forwards to its parent all the messages it collected during the previous rounds. ∎
The following lemma establishes a tight connection between a uniquely-labeled BFS-clustering of a graph $G$ and the virtual graph $H$ resulting from that clustering.
Lemma 7.
Let $G=(V,E)$ be an $n$-node graph, and let $H$ be the virtual graph induced by some uniquely-labeled BFS-clustering $(\ell,\delta)$ of $G$. Let $\mathcal{A}$ be a distributed algorithm running on $H$ with awake complexity $\alpha$ and round complexity $\varrho$. For every vertex $i$ of $H$, let $\textsf{input}(i)$ be the input of vertex $i$ of $H$, and let $\textsf{output}(i)$ be the output of $\mathcal{A}$ at vertex $i$. Assuming each node $v$ of $G$ knows $\textsf{input}(\ell(v))$, it is possible to run $\mathcal{A}$ on $G$ such that every node $v$ of $G$ computes $\textsf{output}(\ell(v))$ with awake complexity at most $7\cdot\alpha$ and round complexity $O(\varrho\cdot n)$.
Proof.
We prove that each round that $\mathcal{A}$ performs on $H$ can be simulated on $G$ by having each node of $G$ awake for at most $7$ rounds during an interval of $R=O(n)$ rounds, where $R$ is the round complexity of computing a convergecast and a broadcast (see Lemma 6). Let $v\in V$ and let $i=\ell(v)$. Node $v$ of $G$ simulates node $i$ of $H$ as follows. Let us call a phase a series of $R$ consecutive rounds. We prove by induction on $k$ that, (1) after $k$ phases (i.e., after $k\cdot R$ rounds), node $v$ is able to compute the state that node $i$ of $H$ would have by executing $k$ rounds of $\mathcal{A}$, and (2) if up to this point node $i$ has been awake for $r$ rounds, then $v$ has been awake for at most $7r$ rounds.
The base case $k=0$ trivially holds. Suppose that the claim holds for all $k^{\prime}$ with $0\leq k^{\prime}<k$, and let us prove the claim for $k$. We distinguish two cases, depending on whether $i$ was awake at round $k$ or not. In the latter case, $v$ does not wake up at all during phase $k$, and the claim trivially holds. In the former case, $v$ wakes up during the first round of phase $k$, and shares its state with all its (awake) neighbors in $G$. If $\delta(v)>0$, then $v$ picks as parent any neighbor $u$ that belongs to the same cluster as $v$, and satisfying $\delta(v)=\delta(u)+1$. Then, $v$ performs a convergecast, followed by a broadcast, as specified in Lemma 6. Hence, in total, $v$ was awake during at most $7$ rounds during phase $i$. During that phase, $v$ learned the state of all the neighbors of $i$ in $H$, and thus it can compute the new state that node $i$ has computed at round $k$. ∎
We now observe that, similarly to what holds in LOCAL model, the execution of two algorithms can be concatenated into one, assuming that all nodes know when the first algorithm terminates.
Lemma 8.
Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be two algorithms with respective awake complexities $S_{1}$ and $S_{2}$, and respective round complexities $T_{1}$ and $T_{2}$. The algorithm obtained by running the two algorithms consecutively, first $\mathcal{A}_{1}$ for $T_{1}$ rounds, and then $\mathcal{A}_{2}$ for $T_{2}$ rounds, has awake complexity $S_{1}+S_{2}$ (and, by design, round complexity $T_{1}+T_{2}$).
Proof.
Nodes first execute $\mathcal{A}_{1}$. When $\mathcal{A}_{1}$ terminates at node $v$, sait at round $t\leq T_{1}$, it becomes idle for $T_{1}-t$ rounds, until round $T_{1}+1$. At round $T_{1}+1$ all nodes start executing $\mathcal{A}_{2}$. For every node $v$, the number of rounds during which $v$ is awake is thus $S_{1}+S_{2}$. ∎
We will use Lemma 8 implicitly throughout the paper, without systematically referring to it.
4 Solving any O-LOCAL Problem Given a Colored BFS-Clustering
In this section, we prove that, if the nodes of a graph $G=(V,E)$ are given a colored BFS-clustering of $G$ as input (i.e., every node knows its color, and its distance value in the clustering), then every problem in the O-LOCAL class can be solved efficiently in $G$. More formally, this section is entirely dedicated to proving the following statement.
Theorem 9.
Let $G=(V,E)$ be an $n$-node graph and let $\Pi$ be a problem in O-LOCAL. Let $(\gamma,\delta)$ be a colored BFS-clustering of $G$, and let us assume that every node $v\in V$ initially knows its color $\gamma(v)$, and its distance $\delta(v)$. Let $c=\max_{v\in V}\gamma(v)$. $\Pi$ can be solved by a distributed algorithm with awake complexity $O(\log c)$, and round complexity $O(c\cdot n)$.
Note that the complexity of the algorithm in the theorem above depends on $\max_{v\in V}\gamma(v)$ but not on $\max_{v\in V}\delta(v)$. The proof of the theorem is based on a result of [BM21] stating that every problem in O-LOCAL has awake complexity $O(\log\Delta+\log^{*}n)$ in the class of graphs with at most $n$ nodes, and maximum degree at most $\Delta$. Let us first briefly provide the reader with a high-level intuition of this result, as it might be useful to the reader for a better understanding of the main ideas allowing us to establish Theorem 9. The algorithm from [BM21] works as follows. First, the nodes compute an $O(\Delta^{2})$-coloring of the graph, in $O(\log^{*}n)$ rounds, by using Linial’s coloring algorithm [Lin92]. Let $q=O(\Delta^{2})$ be the smallest power of $2$ larger than the number of colors produced by Linial’s algorithm. Second, the colors in $\{1,\ldots,q\}$ are mapped to the larger color palette $\{1,\ldots,2q-1\}$. This is done by using a carefully crafted mapping $\phi:\{1,\ldots,q\}\to\{1,\ldots,2q-1\}$ satisfying the following properties.
Lemma 10 (Barenboim and Maimon [BM21]).
There exist two mappings
$\phi:\{1,\ldots,q\}\rightarrow\{1,\ldots,2q-1\},\text{ and }r:\{1,\ldots,q\}\rightarrow 2^{\{1,\ldots,2q-1\}}$
satisfying the following: (1) for every $c\in\{1,\ldots,q\}$, $|r(c)|=1+\log q$, (2) $\phi(c)\in r(c)$, and (3) for every pair of distinct colors $(c_{1},c_{2})\in\{1,\ldots,q\}^{2}$, there exists $x\in r(c_{1})\cap r(c_{2})$ such that
$\min\{\phi(c_{1}),\phi(c_{2})\}<x<\max\{\phi(c_{1}),\phi(c_{2})\}.$
Lemma 10 is proved by considering a binary tree $T$ whose nodes are the elements of the set $\{1,\ldots,2q-1\}$ labeled according to a in-order traversal (see Figure 1). The leaves of $T$ are the images by $\phi$ of the elements in $\{1,\ldots,q\}$, i.e., $\phi(c)$ is the label of the leaf with the $c$-th smallest label among all leaves. For every color $c\in\{1,\ldots,q\}$, $r(c)$ is the set of labels of the nodes in the path from the root of $T$ to the leaf $\phi(c)$.
841226101413579111315
Figure 1: The tree used in the proof of Lemma 10. We have $\phi(2)=3$ as the second smallest label of a leaf is 3, and $r(2)=\{2,3,4,8\}$. Similarly, $\phi(4)=7$, and $r(4)=\{4,6,7,8\}$. Note that the lowest common ancestor of the nodes labeled 3 and 7 is the node labeled 4, and indeed $3<4<7$.
Given the computed $q$-coloring, and given the mapping $\phi$ and $r$, the proof by [BM21] that every problem in O-LOCAL has awake complexity $O(\log\Delta+\log^{*}n)$ continues as follows. Let $G=(V,E)$ be a graph, and let $\mu$ be the orientation of the edges of $G$ resulting from the proper $q$-coloring of $G$, that is, every edge is oriented from its end-point of higher color to its other end-point. For every color $c\in\{1,\dots,q\}$, let us consider the two sets
$r_{<}(c)=\{x\in r(c)\mid x<\phi(c)\},$
and
$r_{>}(c)=\{x\in r(c)\mid x>\phi(c)\}.$
Every node $v\in V$ with color $c=c(v)$ wakes up at every round $i\in r(c)$, and performs the following operations (recall that $\phi(c)\in r(c)$):
  • If $i\in r_{<}(c)$, then node $v$ receives the states of all its neighbors awaken at round $i$, and stores this information in its own state;
  • If $i=\phi(c)$, then node $v$ decides its own solution as a function of its state (which possibly includes the states of some neighbors received during previous rounds);
  • If $i\in r_{>}(c)$, then node $v$ sends its state to its neighbors.
The awake complexity of this algorithm is $O(\log q)$ as, from Lemma 10, $|r(c)|=1+\log(q)$.
To establish that the algorithm correctly produces a solution for the problem $\Pi$, it is sufficient to show that each node $v$ knows the whole graph $G_{\mu}(v)$ when it must compute its local solution to $\Pi$. This can be shown by a simple induction on the colors in $\{1,\dots,q\}$. Indeed, this clearly holds for nodes of color $1$. Now, let us assume that each neighbor $u$ of $v$ with color $c(u)<c(v)$ knows $G_{\mu}(u)$ at round $\phi(c(u))$. By Lemma 10, there exists an element $x\in r(c(u))\cap r(c(v))$ such that $\phi(c(u))<x<\phi(c(v))$. It follows that, at round $x$, node $v$ learns $G_{\mu}(u)$, and thus, node $v$ knows $G_{\mu}(v)$ at round $\phi(c(v))$. We summarize the result of [BM21] by the following lemma for further references in the text.
Lemma 11 (Barenboim and Maimon [BM21]).
[…]
Theorem 9 essentially states that Lemma 11 also holds whenever, instead of a proper coloring of the graph, the algorithm is given a colored BFS-clustering of the graph.
Proof of Theorem 9.
[…]
Claim 12.
[…]
Thanks to Claim 12, we can apply Lemma 11 on $\Pi^{\prime}$. We do so by using the proper coloring $\gamma$ of $H$, and Lemma 11 says that $\Pi^{\prime}$ can be solved in $H$ by a distributed algorithm $A$ with awake complexity $O(\log c)$, and round complexity $O(c)$. We use Lemma 7 for simulating $A$ on $G$ using the uniquely-labeled BFS-clustering $(\ell,\delta)$. Each round of $A$ in $H$ is a virtual round in $G$. Each virtual round is replaced by a sequence of $O(n)$ rounds used for (1) gathering at the root of each cluster $G_{i}$ of $(\ell,\delta)$ the set of all messages exchanged between $G_{i}$ and its neighboring clusters during the previous virtual round, and (2) broadcasting this set to all nodes in $G_{i}$. The total number of rounds consumed for the simulation of $A$ in $G$ is therefore $O(nc)$. Each vertex $i$ of $H$ is awake $O(\log c)$ times during the execution of $A$, and, whenever $i$ is awake, each node of $G_{i}$ is awake a constant number of rounds for performing the convergecast-broadcast operation. It follows that the awake complexity of the simulation of $A$ in $G$ remains constant. This completes the proof of Theorem 9. ∎
5 Computing a Colored BFS-Clustering
In the previous section, we have seen how to solve any problem in O-LOCAL assuming given a colored BFS-clustering. In this section, we show how to compute such a colored BFS-clustering efficiently in the Sleeping model. Specifically, this section is entirely dedicated to establishing the following theorem. Note that Theorem 1, which is the main contribution of the paper, directly follows from combining Theorem 13 with Theorem 9.
Theorem 13.
Let $G=(V,E)$ be an $n$-node graph. A colored BFS-clustering $(\gamma,\delta)$ of $G$ with $\max_{v\in V}\gamma(v)=2^{O(\sqrt{\log n})}$ can be computed by a distributed algorithm with awake complexity $O(\sqrt{\log n}\cdot\log^{*}n)$, and round complexity $O(n^{5}\sqrt{\log n})$.
Remark.
As it will appear clear in the proof, if the IDs assigned to the $n$ nodes are taken from $\{1,\ldots,n\}$, then the round complexity of our algorithm improves to $O(n^{2}\sqrt{\log n})$. More generally, if the IDs are taken from $\{1,\ldots,n^{s}\}$ with $41\leq s<$ 4, then the round complexity of our algorithm becomes $O(n^{1+s}\sqrt{\log n})$.
The rest of the section is entirely dedicated to the proof of Theorem 13. Before entering into the details of the proof, let us establish a couple of statements that will be useful for the proof.
Lemma 14.
Let $G=(V,E)$ be an $n$-node graph, let $H$ be the virtual graph induced by a uniquely-labeled BFS-clustering $(\ell,\delta)$ of $G$, and let $K$ be the virtual graph induced by a uniquely-labeled BFS-clustering $(\ell^{\prime},\delta^{\prime})$ of $H$. There exists a distributed algorithm with constant awake complexity, and round complexity $O(n^{2})$ that computes a uniquely-labeled BFS-clustering (′′$(\ell^{\prime\prime},\delta^{\prime\prime})$ of $G$ such that the virtual graph induced by (′′$(\ell^{\prime\prime},\delta^{\prime\prime})$ is $K$. That is, every node $v\in V$ is given the two pairs $(\ell(v),\delta(v))$ and $(\ell^{\prime}(\ell(v)),\delta^{\prime}(\ell(v)))$ as input, and it must compute the pair (′′$(\ell^{\prime\prime}(v),\delta^{\prime\prime}(v))$.
(a) This figure shows an example of input for Lemma 14. The graph $G$ is depicted in black. The virtual graph $H$ induced by a uniquely-labeled BFS-clustering $(\ell,\delta)$ of $G$ is depicted in orange, where the labels contained in the black nodes denote $\delta$, and $\ell$ is represented by orange circles. The virtual graph $K$ induced by a uniquely-labeled BFS-clustering $(\ell^{\prime},\delta^{\prime})$ of $H$ is depicted in blue, where the labels contained in the orange squares nodes denote $\delta^{\prime}$, and $\ell^{\prime}$ is represented by blue circles.
(b) This figure shows the output of Lemma 14 when applied to the graph in (a). We obtain a uniquely-labeled BFS-clustering (′′$(\ell^{\prime\prime},\delta^{\prime\prime})$ of $G$ such that the virtual graph induced by (′′$(\ell^{\prime\prime},\delta^{\prime\prime})$ is $K$.
Figure 2: An example of application of Lemma 14.
Proof.
A visual representation of Lemma 14 is depicted in Figure 2. Every node $v\in V$ merely sets
′′$\ell^{\prime\prime}(v)=\ell^{\prime}(\ell(v)),$
that is, its cluster is simply resulting from merging the clusters of $(\ell,\delta)$ which are given the same label in $(\ell^{\prime},\delta^{\prime})$. The issue is computing a distance value $\delta^{\prime\prime}(v)$ for $v$. For this purpose, one has first to determine which nodes are roots in the uniquely-labeled BFS-clustering (′′$(\ell^{\prime\prime},\delta^{\prime\prime})$, i.e., which nodes $v$ satisfy $\delta^{\prime\prime}(v)=0$. We set
$\delta^{\prime\prime}(v)=0\iff\delta(v)=0\;\mbox{and}\;\delta^{\prime}(\ell(v))=0.$
That is, a node $v$ with $\ell(v)=i$ is root in (′′$(\ell^{\prime\prime},\delta^{\prime\prime})$ if and only if it is root of its cluster $G_{i}$ in $(\ell,\delta)$, and $G_{i}$ is root in $(\ell^{\prime},\delta^{\prime})$. To compute the value $\delta^{\prime\prime}(v)$ of every node $v$, the nodes of $G$ collectively perform the following operations whose purpose is, for every node $v$, to learn the entire structure of the subgraph of $G$ induced by the set of nodes $\{u\in V\mid\ell^{\prime\prime}(u)=\ell^{\prime\prime}(v)\}=\{u\in V\mid\ell^{\prime}(\ell(u))=\ell^{\prime}(\ell(v))\}$, including the IDs of all its nodes, which is sufficient for computing $\delta^{\prime\prime}(v)$ that is merely equal to the distance of $v$ to the root (of cluster ′′$\ell^{\prime\prime}(v)$) in this subgraph.
First, for each vertex $i$ of $H$, all the nodes $v\in V$ such that $\ell(v)=i$ recover the structure of the subgraph $G_{i}$ of $G$ induced by nodes labeled $i$ by $\ell$. This is merely done by performing in parallel, within each cluster $i$, a convergecast followed by a broadcast scheduled thanks to the function $\delta$. More precisely, every node $v$ with $\ell(v)=i$ sets a message $M_{v}$ containing all edges of $E$ between $v$ and the neighbors $u$ of $v$ in $G$ satisfying $\ell(u)=i$. Moreover, among all these nodes $u$, node $v$ elects as parent $p(v)$ any node satisfying $\delta(u)<\delta(v)$. (If no such node exists, i.e., if $v$ is the root of its cluster, then $v$ sets $p(v)=\bot$.) At this point, we can apply Lemma 6 for allowing the roots, i.e., all nodes $v$ such that $\delta(v)=0$, to acquire the entire structure of their clusters by a convergecast operation (this structure is merely the union of all messages $M_{v}$ received by each root). Then all nodes can acquire the entire structure of their clusters by a broadcast of this structure from the roots within each cluster in parallel. The awake complexity of this computation is constant, and the overall computation lasts $O(n)$ rounds.
To then allow every node $v\in V$ with label ′′$\ell^{\prime\prime}(v)=j$, $j\in V(K)$, to recover the structure of the subgraph $H_{j}=\bigcup_{\ell^{\prime}(i)=j}G_{i}$ of $G$ induced by all nodes $u\in V$ satisfying ′′$\ell^{\prime\prime}(u)=j$, we are first setting a parent function $p^{\prime}$ among the clusters $G_{i}$, $i\in V(H)$, of $(\ell,\delta)$. Every node $v$ set a message $M^{\prime}_{v}$ containing all its incident edges $\{u,v\}$ such that ′′$\ell^{\prime\prime}(u)=\ell^{\prime\prime}(v)$ but $\ell^{\prime}(u)\neq\ell^{\prime}(v)$, plus the distance $\delta^{\prime}(\ell(u))$ for every such edge $\{u,v\}$. By a convergecast operation, the root of every cluster $G_{i}$ can select a neighboring cluster $G_{i^{\prime}}$ with $\ell^{\prime}(i^{\prime})=\ell^{\prime}(i^{\prime})$ and $\delta^{\prime}(i^{\prime})<\delta^{\prime}(i)$, and set $p^{\prime}(i)=i^{\prime}$. It also selects an edge between the two clusters $G_{i}$ and $G_{i^{\prime}}$. All these information are broadcasted to all the nodes in $G_{i}$. Thanks to Lemma 6, the awake complexity of all these operations is constant, and they lasts $O(n)$ rounds.
To allow each node of $H_{j}=\cup_{\ell^{\prime}(i)=j}G_{i}$ to acquire the structure of $H_{j}$, it suffices to simulate in $G$ the execution of a convergecast-broadcast protocol $\mathcal{P}$ in the subgraph of $H$ induced by all vertices $i$ such that $\ell^{\prime}(i)=j$, by applying Lemma 6 using the parent relation $p^{\prime}$ and the distance $\delta^{\prime}$. For this purpose, every virtual round of $\mathcal{P}$ in $H$ is replaced by a sequence of $O(n)$ rounds for allowing all the nodes of every cluster $G_{i}$ of $H_{j}$ to gather all the information exchanged with neighboring clusters during the previous virtual round. Note that, given $n$, $\delta(v)$ and $\delta^{\prime}(\ell(v))$, every vertex can compute the set of rounds during which is must be awake for performing the several convergecast-broadcast operations occurring during the simulation of protocol $\mathcal{P}$.
The total number of rounds required for this simulation is therefore $O(n^{2})$ since each of the $O(n)$ virtual rounds of $\mathcal{P}$ in $H$ requires $O(n)$ rounds to be simulated in $G$. In $\mathcal{P}$, every vertex $i$ of $H$ is activated $O(1)$ times, that is, each cluster $G_{i}$ is awake a constant number of virtual rounds. During a virtual round, each node $v\in V$ is awaken during $O(1)$ rounds. Therefore, every node $v$ is awake a constant number of rounds in total, and thus the total awake complexity of the protocol is constant. ∎
We now state a result that will be the main ingredient for proving Theorem 13. This lemma states that, with awake complexity $O(\log^{*}n)$, one can compute a colored BFS-clustering $(\gamma,\delta)$ of any $n$-node graph $G=(V,E)$ with desirable properties. Namely, the clusters with small colors, i.e., with colors in a range $\{1,\dots,ab^{2}\}$ for well suited parameters $a$ and $b$ of our construction, are reduced to a single node. For clusters with large colors, i.e., colors larger than $ab^{2}$, the colored BFS-clustering is actually a uniquely-labeled BFS-clustering. Moreover, the number of clusters with color $>ab^{2}$ is small, that is, it does not exceed $n/b$ in $n$-node networks. This later property will allow us to proceed by induction in $O(\log_{b}n)$ phases for establishing Theorem 13. For $b=2^{\sqrt{\log n}}$, the number of phases is $O(\log n/\log b)=O(\log n/\sqrt{\log n})=O(\sqrt{\log n})$.
Lemma 15.
There exists an integer $a>0$ such that, for every integer $b>0$, there exists an algorithm parametrized by $b$, with awake complexity $O(\log^{*}n)$ and round complexity $O(n^{4})$ which, given any $n$-node graph $G=(V,E)$, computes a colored BFS-clustering $(\gamma,\delta)$ such that
  • the pair $(\gamma,\delta)$ restricted to the subgraph induced by $\{v\in V\mid\gamma(v)\in\{1,\dots,a\cdot b^{2}\}\}$ is a colored BFS-clustering, and, for every $v\in V$ with $\gamma(v)\in\{1,\dots,a\cdot b^{2}\}$, $\delta(v)=0$ (i.e., $v$ is alone in its cluster);
  • the pair $(\gamma,\delta)$ restricted to the subgraph induced by $\{v\in V\mid\gamma(v)>a\cdot b^{2}\}$ is a uniquely-labeled BFS-clustering with at most $n/b$ clusters.
Let us first show how to prove Theorem 13 assuming that Lemma 15 holds.
Figure 3: Sum up of the Clustering Algorithm of Theorem 13. We compute iteratively a clustering of the remaining clusters. Some parts get their final cluster at each iteration, while the other goes back into the loop.
Proof of Theorem 13.
Let $k=2\sqrt{\log n}$, and $b=2^{\sqrt{\log n}}$. Let us consider the trivial uniquely-labeled BFS-clustering $(\ell_{0},\delta_{0})$ of $G$ obtained by assigning to each node $v$ the label $\ell_{0}(v)=\mbox{ID}(v)$, and $\delta_{0}(v)=0$. Let $H_{0}$ be the virtual graph associated to this clustering. The desired colored BFS-clustering $(\gamma,\delta)$ is computed by a sequence of $k$ phases, labeled by $i=1,\dots,k$ (see Figure 3 for an illustration of the iterative process). Initially, let $G_{0}=G$. For every $i\geq 1$, Phase $i$ consists of the following. We assume given a virtual graph $H_{i-1}$ associated to some uniquely-labeled BFS-clustering $(\ell_{i-1},\delta_{i-1})$ of a subgraph $G_{i-1}=(V_{i-1},E_{i-1})$ of $G$ induced by $V_{i-1}\subseteq V$. That is, every $v\in V_{i-1}$ knows the pair $(\ell_{i-1}(v),\delta_{i-1}(v))$. Phase $i$ consists of the following.
Thanks to Lemma 15 applied for $b=2^{\sqrt{\log n}}$, a colored BFS-clustering $(\gamma^{\prime}_{i-1},\delta^{\prime}_{i-1})$ of $H_{i-1}$ can be computed, with awake complexity $O(\log^{*}n)$, and round complexity $O(n^{4})$ on the virtual graph $H_{i-1}$. Let
$S_{i-1}=\{C\in V(H_{i-1})\mid\gamma^{\prime}_{i-1}(C)\in\{1,\dots,ab^{2}\}\}.$
and
$L_{i-1}=\{C\in V(H_{i-1})\mid\gamma^{\prime}_{i-1}(C)>ab^{2}\},$
The colored BFS-clustering $(\gamma^{\prime}_{i-1},\delta^{\prime}_{i-1})$ of $H_{i-1}$ has the property that, for each vertex $C\in S_{i-1}$, the cluster of $C$ consists of $C$ only. On the other hand, restricted to the vertices $C\in L_{i-1}$, $(\gamma^{\prime}_{i-1},\delta^{\prime}_{i-1})$ is a uniquely-labeled BFS-clustering. Our aim is to simulate on $G_{i-1}$ the computation of $(\gamma^{\prime}_{i-1},\delta^{\prime}_{i-1})$ in $H_{i-1}$ so that every node $v\in V_{i-1}$ computes the pair
$(\gamma^{\prime}_{i-1}(\ell_{i-1}(v)),\delta^{\prime}_{i-1}(\ell_{i-1}(v))).$
This can be achieved by a mere application of Lemma 7, with awake complexity $O(\log^{*}n)$, and round complexity $O(n^{5})$. At this point, every node $v\in V_{i-1}$ satisfying $\gamma^{\prime}_{i-1}(\ell_{i-1}(v))\in\{1,\dots,ab^{2}\}$, sets
$\gamma(v)=(i,\gamma^{\prime}_{i-1}(\ell_{i-1}(v))),\;\mbox{and}\;\delta(v)=\delta_{i-1}(v),$
and terminates. For all the other nodes, i.e., for all the nodes of $G_{i-1}$ whose clusters in $(\gamma^{\prime}_{i-1},\delta^{\prime}_{i-1})$ belong to $L_{i-1}$, one can apply Lemma 14, stating that the nodes of $G_{i-1}$ can compute a uniquely-labeled BFS-clustering $(\ell_{i},\delta_{i})$ such that, for every $v\in V_{i-1}$,
$\ell_{i}(v)=\gamma^{\prime}_{i-1}(\ell_{i-1}(v)),$
and $\delta_{i}(v)$ is its distance to the root of its cluster. The graph $G_{i}$ is the subgraph of $G$ induced by $V_{i}=\{v\in V_{i-1}\mid\ell_{i}(v)\text{ is defined and }\ell_{i}(v)>ab^{2}\}$. The graph $H_{i}$ is the virtual graph of the uniquely-labeled BFS-clustering $(\ell_{i},\delta_{i})$ of $G_{i}$. Note that, thanks to Lemma 15, $|V(H_{i})|\leq|V(H_{i-1})|/b$.
The total awake complexity of the $k$ iterations is $O(\sqrt{\log n}\cdot\log^{*}n)$, and these $k$ iterations consume $O(n^{5}\sqrt{\log n})$ rounds in total. After $k$ iterations of the above sequence of instructions, the virtual graph $H_{k}$ is empty. Indeed, the number of remaining nodes after $k$ iterations can be upper bounded by
$\frac{n}{b^{k}}=\frac{n}{2^{2\sqrt{\log n}\cdot\sqrt{\log n}}}=\frac{n}{2^{2\cdot\log n}}=\frac{1}{n}<1.$
It follows that, after $k$ iterations, all nodes of $G$ have terminated. The number of colors used by our construction can be upper bounded by
$k\cdot a\cdot b^{2}=2^{O(\sqrt{\log n})},$
as desired. It remains to show correctness, i.e., that $(\gamma,\delta)$ is indeed a colored BFS-clustering of $G$.
By construction, nodes assigned to clusters at different iterations have different colors, and therefore two clusters created at different iterations have different colors. Let $C\subseteq V$ be a cluster created at some iteration $i\in\{1,\dots,k\}$. Every node $v\in C$ of this cluster is labeled $\gamma(v)=(i,\gamma^{\prime}_{i-1}(\ell_{i-1}(v)))$. The cluster $C$ corresponds to vertex $C\in V(H_{i-1})$ with $\gamma^{\prime}_{i-1}(C)\in\{1,\dots,ab^{2}\}$, where $H_{i-1}$ is the virtual graph corresponding to the uniquely-labeled BFS-clustering $(\ell_{i-1},\delta_{i-1})$ of $G_{i-1}$. In the clustering $(\gamma^{\prime}_{i-1},\delta^{\prime}_{i-1})$, $C$ is its own cluster. It follows that the setting of $\delta(v)=\delta_{i-1}(v)$ satisfies the property of a BFS-clustering.
Finally, let us consider two nodes $v$ and $v^{\prime}$ of $G$ belonging to clusters $C$ and $C^{\prime}$ of $\gamma$, with
$\gamma(v)=(i,\gamma^{\prime}_{i-1}(\ell_{i-1}(v)))=\gamma(v^{\prime})=(i,\gamma^{\prime}_{i-1}(\ell_{i-1}(v^{\prime}))).$
In this case, $C$ and $C^{\prime}$ were clusters of $G_{i-1}$, and, in the clustering $\gamma^{\prime}_{i-1}$, $C$ and $C^{\prime}$ are singletons. If there is an edge between $v$ and $v^{\prime}$ in $G$, then since $(\ell_{i-1},\delta_{i-1})$ is a uniquely-labeled BFS-clustering of $G_{i-1}$, we have that $C=C^{\prime}$. It follows that $(\gamma,\delta)$ is a colored BFS-clustering, as desired. ∎
For completing the proof of Theorem 13, it is thus sufficient to prove Lemma 15. That is, we will provide an algorithm that computes a uniquely-labeled BFS-clustering satisfying the requirements of Lemma 15, with awake complexity $O(\log^{*}n)$, and round complexity $O(n^{4})$.
Proof of Lemma 15.
13181451651201361131630105102142113101121135145126133157128
(a) Parent selection
0262626363737272636372023302032042420273232
(b) Cluster decomposition
Figure 4: In (a), we have a distance-2 coloring $c_{1}$ of the nodes, with $b=3$ and $k=100$. Notice that nodes of degree $\leq 3$ have added 100 to their colors. In red, each node $u$ is connected to $p_{1}(u)$ (with a self loop if $p_{1}(u)=\bot$). In (b), each node $u$ has computed its new color $c_{2}(u)$ (that is not a proper coloring, it is a decreasing coloring from node to parent) and its new parent $p_{2}(u)$ (in blue if it differs from $p_{1}(u)$). The trees formed with a root of degree at least 4 form the new clusters. (Grey) Nodes in a tree of root of degree at most 3 become singleton clusters (dotted tree edges to illustrate that the parent relation is forgotten). These nodes will compute some $ab^{2}$ coloring.
The algorithm works as follows (see Figure 4). First, the nodes of the $n$-node graph $G$ collectively compute a proper coloring of $G^{2}$ with a palette of $k=O(n^{4})$ colors, where $G^{2}$ is the square of $G$, that is, the graph with the same set of nodes as $G$, and where two nodes are connected by an edge if they are at distance at most $2$ in $G$. This can be done in $O(\log^{*}n)$ rounds in the LOCAL model by using, e.g., Linial’s coloring algorithm [Lin92]. Let us denote by $c_{0}:V\to\{1,\dots,k\}$ this coloring.
Note that if the IDs assigned to the nodes are taken from the range $\{1,\ldots,n^{s}\}$ for some $s\geq 1$, then this step could be replaced by directly taking the IDs of the nodes as the desired distance-$2$ $k$-coloring, with $k=n^{s}$, in zero rounds.
Let $b>0$ be an integer. Every node $v$ with $\deg_{G}(v)\leq b$ increases its color $c_{0}(v)$ by adding $k$ to it. That is, the new color $c_{1}(v)$ of $v$ is set as $c_{1}(v)=c_{0}(v)+k$. Every node $v$ with $\deg_{G}(v)>b$ sets $c_{1}(v)=c_{0}(v)$. Hence, for every $v\in V$, we have $1\leq c_{1}(v)\leq 2k$. Note that the colors of the nodes of degree at most $b$ are in the range $\{k+1,\dots,2k\}$, whereas the colors of the nodes of degree larger than $b$ are in the range $\{1,\dots,k\}$.
Let us now set a “parent pointer” $p_{1}(v)$ at every node $v\in V$. Let us denote by $N(v)$ the (open) neighborhood of $v$, and by $N^{2}(v)$ the set of nodes at distance exactly 2 from $v$ in $G$. Three cases are considered, depending on the colors of the nodes in $N(v)\cup N^{2}(v)$.
  • If every node $u\in N(v)\cup N^{2}(v)$ satisfies $c_{1}(u)>c_{1}(v)$, then $p_{1}(v)=\bot$. A “shift” variable $b(v)$ is also set to $b(v)=\bot$ in this case.
  • If there exists $u\in N(v)$ satisfying $c_{1}(u)<c_{1}(v)$, then $p_{1}(v)=u_{min}$ where $u_{min}$ is the node with smallest color in $N(v)$, breaking ties arbitrarily. In this case, the shift at $v$ is set to $b(v)=0$.
  • If none of the above two conditions are fulfilled, that is, if every $u\in N(v)$ satisfies $c_{1}(u)>c_{1}(v)$, but there exists $u\in N^{2}(v)$ such that $c_{1}(u)<c_{1}(v)$, then $p_{1}(v)$ is set as the node with smallest color in $N^{2}(v)$, breaking ties arbitrarily. In this case, the shift at $v$ is set to $b(v)=1$.
By construction, the collection of pointers $p_{1}$ induces a rooted spanning forest $F_{1}$ of $G^{2}$ (i.e., every tree in the forest is rooted). Namely, the roots are all nodes $v$ with $p_{1}(v)=\bot$. Moreover, for every $v\in V$,
$p_{1}(v)\neq\bot\Longrightarrow c_{1}(v)>c_{1}(p(v)).$
The pointers $p_{1}$, and the coloring $c_{1}$ are now modified at every node to obtain a new pointer $p_{2}$, and a new coloring $c_{2}$, as follows:
  • Every root $v$ of a tree in $F_{1}$, i.e., every node $v\in V$ with $p_{1}(v)=\bot$, sets its color $c_{2}(v)=0$. Every other node $v$ takes color
    $c_{2}(v)=2\cdot c_{1}(p_{1}(v))+b(v).$
    That is, every $v\in V$ with $p_{1}(v)\neq\bot$ copies the color of its parent, doubles it, and shifts it by $1$ if its parent is at distance $2$ from it in $G$.
  • The new pointer $p_{2}(v)$ at node $v\in V$ is defined as follows:
    • If $p_{1}(v)=\bot$, then $p_{2}(v)=\bot$;
    • If $b(v)=0$, then $p_{2}(v)=p_{1}(v)$;
    • If $b(v)=1$, then $p_{2}(v)=u$ where $u$ is an arbitrary node in $N(v)\cap N(p_{1}(v))$.
    Note that we have $p_{2}(v)=\bot$ if and only if $p_{1}(v)=\bot$.
Claim 16.
For every node $v\in V$ with $p_{2}(v)\neq\bot$, $c_{2}(v)>c_{2}(p_{2}(v))$, and the collection of pointers $p_{2}$ induces a rooted spanning forest $F_{2}$ in $G$.
This claim will be proved later, and we carry on with the description of the algorithm by assuming that it holds.
In each rooted tree in $F_{2}$, the nodes collectively perform a convergecast, followed by a broadcast, in order to acquire the whole structure of the tree, including the IDs and colors $c_{2}$ of all its nodes. In particular, every node $v$ can compute its distance $\delta_{aux}(v)$ from the root of its tree, and the ID of this root, that we denote by $\ell(v)$. By construction, the pair $(\ell,\delta_{aux})$ is a uniquely-labeled BFS-clustering of $G$. Note that, thanks to Lemma 6, the convergecast-broadcast operation can be done in $O(n^{4})$ rounds, with constant awake complexity. The colored BFS-clustering $(\gamma,\delta)$ of Lemma 15 is obtained from $(\ell,\delta_{aux})$ as follows.
Let $C$ be a cluster in $(\ell,\delta_{aux})$, with a root $r$ satisfying $\deg(r)\leq b$. It must be the case that every node $v\in C\smallsetminus\{r\}$ satisfies $c_{2}(v)>c_{2}(r)$. Moreover, since the colors of the nodes with degree $\leq b$ are larger than the colors of the nodes with degree $>b$, we get that, for every $v\in C$, $\deg(v)\leq b$. Let $U\subseteq V$ be the subset of nodes in $G$ belonging to a cluster $C$ whose root $r$ satisfies $\deg(r)\leq b$, and let $G[U]$ denote the subgraph of $G$ induced by the nodes in $U$. By construction, the graph $G[U]$ has maximum degree at most $b$. The nodes of $G[U]$ collectively compute a proper coloring $\gamma$ of $G[U]$ with $O(b^{2})$ colors using Linial’s coloring algorithm [Lin92]. Let us fix $a$ as the smallest integer such that the number of colors produced by Linial’s algorithm is at most $a\cdot b^{2}$. The round complexity (and therefore the awake complexity) of this algorithm is $O(\log^{*}n)$. At this point, every $v\in U$ becomes part of a cluster $\gamma(v)$ consisting of $v$ alone, and $v$ sets $\delta(v)=0$.
Every node $v\in V\smallsetminus U$, i.e., every node $v$ belonging to a cluster in $(\ell,\delta_{aux})$ with a root of degree at least $b+1$ sets $\gamma(v)=\ell(v)+a\cdot b^{2}$, and $\delta(v)=\delta_{aux}(v)$.
This completes the construction of the desired clustering $(\gamma,\delta)$. The total awake complexity of all the above operations is $O(\log^{*}n)$, and its round complexity is $O(n^{4})$. It remains to show correctness (including the proof of Claim 16).
By construction, the colored BFS-clustering $(\gamma,\delta)$ satisfies the statement of Lemma 15 regarding the nodes in $U$ and the nodes in $V\smallsetminus U$. What is left to prove is that, if we restrict the BFS-clustering $(\gamma,\delta)$ to the subgraph induced by $\{v\in V\mid\gamma(v)>a\cdot b^{2}\}$, we obtain at most $n/b$ clusters. For this purpose, we upper bound the number of roots. For each cluster $C$ in $(\ell,\delta_{aux})$, the root $r$ of $C$ satisfies $\deg(r)>b$. Moreover, for each neighbor $v$ of $r$, $v$ cannot be neighbor of a root $r^{\prime}$ of another cluster $C^{\prime}$ in $(\ell,\delta_{aux})$ with $\deg(r^{\prime})>b$. This is because the roots in $(\ell,\delta_{aux})$ are local minima in $G^{2}$, and thus they are at mutual distance at least $3$ in $G$. It follows that, for each root $r$ in $(\ell,\delta_{aux})$ with $\deg(r)>b$, we can charge all $r$’s neighbors to $r$. These neighbors are at least $b+1$, and they are not roots. Note that each node $v\in V$ is charged to at most one root. We thus get that the number of roots $r$ with $\deg(r)>b$ are at most $n/b$, as desired. This completes the proof of Lemma 15, assuming Claim 16 holds. ∎
It just remains to prove Claim 16.
Proof of Claim 16.
We first show that, for every node $v\in V$ with $p_{2}(v)\neq\bot$, $c_{2}(v)>c_{2}(p_{2}(v))$. For each such node $v$, we have $b(v)\in\{0,1\}$. We analyse the two cases separately.
  • Let $v\in V$ with $b(v)=0$, i.e., $p_{2}(v)=p_{1}(v)$. In this case, node $p_{1}(v)$ may either be a root or not. If $p_{1}(p_{1}(v))=\bot$, then
    $c_{2}(p_{2}(v))=c_{2}(p_{1}(v))=0,\;\mbox{and}\;c_{2}(v)=2\cdot c_{1}(p_{1}(v))>0,$
    from which it follows that $c_{2}(p_{2}(v))<c_{2}(v)$. If $p_{1}(p_{1}(v))\neq\bot$, say $p_{1}(p_{1}(v))=u$, then $c_{1}(p_{1}(v))>c_{1}(u)$. Moreover,
    $c_{2}(p_{2}(v))=c_{2}(p_{1}(v))=2\cdot c_{1}(u),\;\mbox{and}\;c_{2}(v)=2\cdot c_{1}(p_{1}(v)),$
    from which it again follows that $c_{2}(p_{2}(v))<c_{2}(v)$.
  • Let $v\in V$ with $b(v)=1$. Let $p_{2}(v)=u\in N(v)\cap N(p_{1}(v))$. By the definition of $b(v)$, $c_{1}(u)>c_{1}(v)$, and $b(u)=0$. Since $u$ is a neighbor of $p_{1}(v)$, $p_{1}(v)\in N(u)$, and since $p_{1}(v)\in N^{2}(v)$, we get that $p_{1}(v)$ has the smallest color among all nodes in $N(u)$. Moreover, since $c_{1}$ is a distance-2 coloring, there are no nodes $z\in N(u)\smallsetminus\{p_{1}(v)\}$ that have the same color as $p_{1}(v)$. By combining these facts with the fact that $b(u)=0$, we get that $p_{1}(u)=p_{1}(v)$. As a consequence,
    $c_{2}(v)=2\cdot c_{1}(p_{1}(v))+1>2\cdot c_{1}(p_{1}(v))=2\cdot c_{1}(p_{1}(u))=c_{2}(u)=c_{2}(p_{2}(v)).$
So the first statement of Claim 16 holds in the two cases $b(v)=0$ or $1$.
Finally, to show that the collection of pointers $p_{2}$ induces a spanning rooted forest $F_{2}$ in $G$, it is sufficient to note that each node $v\in V$ with $p_{2}(v)\neq\bot$ satisfies $c_{2}(v)>c_{2}(p_{2}(v))$. It follows that each such node $v$ is part of a rooted tree, and the union of all these trees spans $G$. This completes the proof of Claim 16, and thus the proof of Lemma 15. ∎
6 Conclusion and Open Problems
We have proved that the awake complexity of computing a colored BFS-clustering with $2^{O(\sqrt{\log n})}$ colors is $O(\sqrt{\log n}\log^{*}n)$ rounds. Moreover, we can exploit such a clustering to solve any problem in the O-LOCAL class with $O(\sqrt{\log n})$ awake complexity. We do not know whether the $\log^{*}n$ term in the awake complexity of computing this clustering is necessary.
Open Question 1.
Is it possible to compute a colored BFS-clustering with $2^{O(\sqrt{\log n})}$ colors with awake complexity $O(\sqrt{\log n})$?
More importantly, it would be interesting to figure out whether the $\sqrt{\log n}$ bound can be reduced for all problems in the O-LOCAL class.
Open Question 2.
Is it possible to solve all problems in O-LOCAL with $o(\sqrt{\log n})$ deterministic awake complexity?
Additionally, while we focused mostly on the awake complexity of our algorithms, their round complexities are polynomial. It would thus be interesting to figure out whether similar results as the one in the paper regarding the awake complexity can ba achieved while improving the round complexity to, e.g., polylogarithmic. Similarly, it would be interesting to decrease the average awake complexity of the nodes.
Open Question 3.
Is it possible to solve all problems in O-LOCAL with $o(\log n)$ deterministic awake complexity and polylogarithmic round complexity? Is it possible to do so with $o(\sqrt{\log n})$, or even constant average awake complexity?
Finally, understanding how much randomization can help is also an interesting open question.
Open Question 4.
Can we exploit randomization for solving problems in O-LOCAL with $o(\sqrt{\log n})$ awake complexity?